### v2

# FoMo-Bench: a multi-modal, multi-scale and multi-task Forest Monitoring Benchmark for remote sensing foundation models

Abstract
Forests are an essential part of Earth’s ecosystems and natural systems, as well as providing services on which humanity depends, yet they are rapidly changing as a result of land use decisions and climate change. Understanding and mitigating negative effects requires parsing data on forests at global scale from a broad array of sensory modalities, and recently many such problems have been approached using machine learning algorithms for remote sensing. To date, forest-monitoring problems have largely been addressed in isolation. Inspired by the rise of foundation models for computer vision and remote sensing, we here present the first unified Forest Monitoring Benchmark (FoMo-Bench). FoMo-Bench consists of 15 diverse datasets encompassing satellite, aerial, and inventory data, covering a variety of geographical regions, and including multispectral, red-green-blue, synthetic aperture radar (SAR) and LiDAR data with various temporal, spatial and spectral resolutions. FoMo-Bench includes multiple types of forest-monitoring tasks, spanning classification, segmentation, and object detection. To further enhance the diversity of tasks and geographies represented in FoMo-Bench, we introduce a novel global dataset, TalloS, combining satellite imagery with ground-based annotations for tree species classification, encompassing 1,000+ categories across multiple hierarchical taxonomic levels (species, genus, family). Finally, we propose FoMo-Net, a baseline foundation model with the capacity to process any combination of commonly used spectral bands in remote sensing, across diverse ground sampling distances and geographical locations worldwide. This work aims to inspire research collaborations between machine learning and forest biology researchers in exploring scalable multi-modal and multi-task models for forest monitoring. All code and data will be made publicly at https://github.com/RolnickLab/FoMo-Bench.
Keywords:
Forest monitoring Land use Remote sensing AI for Good Multi-sensor data Dataset1 Introduction
Forests worldwide are under threat from climate change, land use change, and invasive species [4, 10, 24, 63], leading to significant biodiversity decline and economic losses. However, effective forest management and conservation are hampered by a lack of information, highlighting the need for scalable forest monitoring. This gap includes multiple specific problems, such as tree species identification, tree size assessment, land cover classification, etc., which also vary considerably across geographies.
Remote sensing has been explored for forest monitoring using various sensors and spatial scales [22, 43], and increasingly, deep learning methods have been used to address various specific forest-related problems [34]. Simultaneously, the field of computer vision has seen a growth in foundation models [3], i.e. models flexible to diverse modalities, scales and data regimes while performing well across downstream tasks [36, 31, 30]. However, it has been noted that general-purpose foundation models in computer vision may be ineffective in specific problem domains [55], presenting a need for foundation models explicitly designed for specific problem spaces, such as Earth observation [38] and weather or climate modeling [44].
In this work, we propose a framework to evaluate foundation models for forest-monitoring tasks. Our Forest Monitoring Benchmark (FoMo-Bench) brings together a wide range of publicly available datasets for forest monitoring, encompassing significant spatial, temporal, and spectral diversity as illustrated in Figure 1. The benchmark includes a mixture of classification, segmentation and detection tasks, from land cover classification to tree crown segmentation. To improve the geographic diversity of FoMo-Bench, we also propose TalloS, a unique, global, multi-modal dataset focused on tree species identification and based on worldwide forest inventories [32]. Each sample included in TalloS combines data acquired from the Sentinel-2 and Sentinel-1 missions, meteorological data from the ERA-5 dataset and elevation information, and georeferenced tree species information, in the form of a multi-label classification task. Finally, we use FoMo-Bench combined with unlabeled data archives to develop a baseline for multi-modal foundation models, namely FoMo-Net, which can flexibly process most commonly used spectral bands in remote sensing. Despite being constrained to a frozen encoder setting, FoMo-Net competes with supervised baselines for semantic segmentation tasks, though it underperforms supervised approaches on multi-label classification and object detection tasks. FoMo-Net acts as a baseline for an ambitious research question: Can we unify foundation models trained on specific individual remote sensors and ground sampling distances (GSD) into a singular, versatile foundation model applicable and adaptable to the requirements (e.g. combination of spectral bands, GSDs, and geographic location) of a wide range of downstream tasks? Overall, our contributions can be summarized as follows:
We provide the first, curated benchmark for forest monitoring, with a global spatial coverage and a diverse set of evaluation tasks;
We introduce a novel multi-modal dataset for tree species multi-label classification;
We introduce FoMo-Net, a unique pre-training approach designed to train foundation models with great flexibilty in terms of the input bands, their ground sampling distance, and geographic location.
2 Related Work
Large-scale remote sensing data has been available for many years through state-funded projects such as Sentinel and Landsat. Building on such sources, various machine learning benchmarks have been developed, such as BigEarthNet [59], Sen12MS [57], Sen4AgriNet [61] and CropHarvest [66]. These initiatives have facilitated extensive research, particularly in land use and land cover (LULC) estimation and forest-related tasks such as tree species identification [14].
Recent exploration of deep learning methods for forest monitoring [34] has been made possible by a range of specific datasets at various scales and geographic locations [46]. High resolution image recordings from unmanned aerial vehicles (UAVs) are part of recent datasets for automated canopy mapping [17], woody invasive species detection [35] and tree species detection [33]. In particular, ReforesTree [53] combines UAV recordings with inventories for both detecting trees and estimate their aboveground biomass. The NeonTree dataset [68] is unique in grouping airborne recordings including light detection and ranging (LiDAR), red-blue-green (RGB) and hyperspectral measurements. Data recorded from satellite sensors have been used for deforestation detection [29] and canopy height estimation [48] using Landsat, or vegetation cover estimation using Sentinel [57]. Although inventories have been used alongside satellite data in Russia with less than a thousand samples [6, 21], we are not aware of any remote sensing dataset providing fine annotations on a large volumetric and geographical scale. Recently, manual forest inventories, including manual measurements of tree height and aboveground biomass, have been aggregated in Tallo [32], a data source that we leverage for machine learning in the present work.
Increasingly, benchmarks across the space of computer vision have moved beyond single tasks to encompass a diversity of related tasks. The large scale BigEarthNet-MM dataset [59], built from Sentinel-1 and Sentinel-2 data, is designed to assess both performance and training complexity of deep learning methods [47]. The SustainBench [70] combines various satellite imagery and street-level images to monitor sustainable development goals covering 15 different topics. Geo-Bench [37] proposes to group remote sensing datasets – including SAR, multispectral, elevation and cloud probabilities data – to tackle six classification and six segmentation tasks focused on LULC. However, the spatial coverage of Geo-Bench is narrow, reducing its potential for global applicability. There remains no benchmark focused on forest monitoring worldwide, combining the full suite of forest inventories, aerial and satellite data, and including the several relevant classification, segmentation and detection tasks.
Principally driven by LULC, foundation models have been investigated for multi-modal and multi-task remote sensing [60, 8] as well as focused on specific satellite products. Recent methods have exploited cropped satellite imagery to simulate various GSDs e.g. [52], as well as contrastive losses e.g. CROMA and Cross-Scale MAE [16, 62]. These methods however lack diversity in the supported spectral bands and GSDs. The recently introduced Presto algorithm [67] is able to handle a variable number of spectral bands, but is restricted to single-pixel timeseries. The SpectralGPT approach [28], leverages the diversity of multispectral bands through the use of 3D tokens, which presents scalability challenges when applied to a vast number of bands in practice. Furthermore, while SkySense [23] stands as the largest and high-performing remote sensing foundational model according to existing benchmarks to date, its architecture incorporates a distinct encoder for each group of spectral bands, which limits its adaptability to new spectral band sets introduced by novel sensors. To the best of our knowledge, foundation models have not been yet applied to multiple forest monitoring tasks at multiple resolutions (a few centimeters to tens of meters) while including a notable diversity of spectral bands, despite significant potential identified in this area [46].
3 FoMo-Bench
There is a wide variety of different open-access forest monitoring datasets, with great diversity in terms of sensors as well as spatial, spectral and temporal resolutions [46], and encompassing a variety of tasks from LULC classification to tree species identification. With the multitude of potential forest monitoring tasks, curating datasets for each is impractical. Foundation models offer the potential for a flexible backbone that can rapidly be fine-tuned on smaller annotated datasets. We here introduce FoMo-Bench, which offers a robust testbed for forest-monitoring foundation models. By carefully selecting datasets composing FoMo-Bench, we have ensured that it (1) encompasses a diversity of geographic locations and environmental conditions, (2) spans a range of task types, and (3) includes input data from a wide variety of sensors e.g. optical, SAR, multispectral, DEM and LiDAR, as well as a range of ground sampling distances, from a few centimeters to per pixel. The diversity of the FoMo-Bench data and tasks is illustrated in Figure 1 and described in more detail below. We have curated all the datasets in FoMo-Bench, spatially split them to train machine learning models and created a reproducible environment for future work with a unified set of PyTorch data loaders.
Satellite-based datasets: FoMo-Bench includes seven datasets involving satellite imagery, namely BigEarthNet-MM, Sen12MS, RapidAI4EO, ForestNet, FiveBillionPixels, TreeSatAI, and TalloS. The BigEarthNet-MM [59] and Sen12MS [57] datasets are two of the most commonly used multi-modal datasets for LULC in remote sensing. Both are based on the Copernicus Sentinel missions offering SAR and multispectral data from Sentinel-1 and Sentinel-2 respectively. For the BigEarthNet-MM dataset, we adapt the existing label structure (based on the Corine land cover database [12]) to encompass 19 distinct land cover classes including broad-leaved, coniferous and mixed forests. While BigEarthNet-MM covers 10 countries in Europe, the Sen12MS dataset has a global spatial coverage (see Figure 2(a)), offering annotations for 17 classes based on MODIS land cover maps [15] with 5 different classes for forest cover. FoMo-Bench also includes RapidAI4EO [2] as a third LULC dataset with a unique source of high resolution satellite imagery offering timeseries of Planet data, at 3m resolution, for 500,000 locations across Europe. We adapt RapidAI4EO labels to encompass 24 classes, 12 of them relevant to forest monitoring. The ForestNet dataset [29] aims to identify drivers leading to deforestation and uses Landsat-8 imagery. FiveBillionPixels dataset [64] provides high resolution Gaofen-2 images () for LULC of cities in China with 24 categories including 5 forest monitoring classes. The TreeSatAI dataset [1] offers multi-modal data (e.g. aerial images, SAR and multispectral) for tree species identification, including 20 tree species derived from forest administration data in Lower Saxony, Germany. Finally, we introduce a completely novel dataset, TalloS, detailed in Section 4.
Aerial-based datasets: The limited spatial resolution offered by free satellite data is not suited for certain remote sensing tasks, such as tree canopy detection or segmentation. FoMo-Bench includes eight datasets with high resolution data from aerial sensor data, namely NeonTree, Woody, ReforesTree, Spekboom, Waititu, FLAIR #1, FLAIR #2 and FORinstance. The NeonTree dataset [68] focuses on tree crown detection with RGB imagery, LiDAR and hyperspectral data recorded from airborne systems. FoMo-Bench includes the NeonTree high resolution RGB imagery (0.1m resolution) along with point cloud LiDAR data. The Woody dataset [35] is composed of RGB images recorded from an UAV, at several cm resolution, for detection of several invasive tree species in Chile. Similarly, the Spekboom [17] and Waititu [33] datasets are composed of RGB images recorded from UAVs to map tree species in South Africa and New Zealand, respectively. In the same vein, the ReforesTree dataset [53] is designed for tree crown detection with 6 classes using high-resolution RGB images recorded from an UAV, as well as ground measurements, such as tree height and diameter at breast height. The FORinstance dataset [49] provides a unique source of data with LiDAR point clouds and semantic annotations collected in 5 countries (see Fig. 2(a)). Unlike the previous aerial datasets, FLAIR #1 [20, 19] and FLAIR #2 [18, 19] are large-scale datasets testing spatiotemporal generalization. Both datasets cover wide areas in France, providing high resolution aerial images (0.2m) and elevation maps (plus Sentinel-2 timeseries in the case of FLAIR #2). These datasets test LULC classification according to 19 classes, including coniferous and deciduous forest.
4 The TalloS dataset
| Split | Number of Classes | Number of Samples |
| Train | 1159 | 161,332 |
| Val | 24 | 4,049 |
| Test | 223 | 25,100 |
| Total | 1160 | 190,481 |
As shown in Fig. 2, the spatial coverage of pre-existing datasets in FoMo-Bench is limited. We improve the diversity and spatial coverage of FoMo-Bench by designing a new dataset, TalloS. This dataset couples manual tree inventories with multispectral and SAR timeseries imagery for a fine grained multi-label classification challenge.
As described in the previous section, various tasks in forest monitoring have required high-resolution data from UAVs but these typically have limited spatial coverage due to human and hardware limits [17, 53, 35, 33]. Satellite-based datasets with fine-grained annotations have previously been lacking for forest monitoring applications.
TalloS is a large-scale multi-modal dataset of satellite-based imagery for tree species multi-label classification. It is based on the global Tallo database [32], containing almost manual georeferenced recordings aggregating information from forest inventories including the tree species as well as tree height, stem diameter, and crown radius. Due to the modest ground sampling distance of publicly accessible satellites (), we design TalloS to concentrate on areas containing a minimum of 50 trees, as regions with lower tree populations lack distinctive signatures in satellite imagery.
TalloS data represent a time series from 2016 to 2023 with a temporal frequency of 15 days. Each timestep consists of a multispectral image from Sentinel-2, a SAR image from Sentinel-1, a digital elevation model (DEM) from Copernicus GLO-30, and meteorological information from ERA5 [27]. Each remote sensing sample of TalloS is mapped to ground-validated and fine grained measurements of trees in its given area, including tree species, genus and family as well as their specific traits. For the purpose of this study, we concentrate on the genus level labels, as distinguishing between species may require higher resolution imagery. Throughout this paper, unless explicitly stated otherwise, we refer to the various genus categories of TalloS as its classes. As shown in Table 1 and Figure 3, TalloS represents a fine-grained classification problem involving more than 1,000 classes with a heavily skewed data distribution. The task is framed as multi-label classification, with the goal of predicting every tree genus that is depicted in the image timeseries.
In Figure 2(b), we illustrate the wide spatial distribution of TalloS as well as the proposed training, validation and test splits. Our goal was to maximize the geographic spatial coverage of each split while ensuring (a) that regions used for training and testing did not overlap, which could lead to inaccurate evaluation, (b) that most classes present in the validation and test sets are also included in the training set. Accommodating both conditions proved challenging due to the limited geographic distribution of many tree genera. The final dataset split is shown in Table 1.
5 FoMo-Net
Building on FoMo-Bench and inspired by the necessity for a generalized foundation model, we introduce FoMo-Net. FoMo-Net is a first attempt to process the most common modalities in the remote sensing domain with a single, sensor-agnostic foundation model, offering a robust foundation model baseline for FoMo-Bench.
Our approach is driven by three core considerations. First, it should maximize flexibility to process various input settings, without relying on specific sets of sensory modalities. Second, it should have the capacity to process information and generate meaningful representations for the whole globe. Third, it should be applicable to a wide range of downstream tasks.
Pre-training data: FoMo-Net pre-training scheme contains rich multi-sensor information, both paired and unpaired, from most parts of the world. In particular, we use the RapidAI4EO dataset as a source of Planet and Sentinel-2 data; TalloS containing DEM, Sentinel-1, and Sentinel-2; SSL4EO-Landsat [58] to acquire global information from Landsat 8 and 9; a combination of all available UAV datasets in FoMo-Bench; and the FiveBillionPixel dataset providing a unique source of Gaofen-2 high resolution satellite imagery.
| Dataset | Sensors | Spatial resolution | Sampling weight |
| SSL4EO-Landsat | Landsat 8-9 | Low | 0.2 |
| RapidAI4EO | Planet, Sentinel-2 | Medium, Low | 0.2 |
| TalloS | Sentinel-1, Sentinel-2, DEM | Low | 0.2 |
| FLAIR #1 | Aerial | High | 0.1 |
| FiveBillionPixels | Gaofen-2 | Medium | 0.2 |
| UAV-datasets | Aerial | Very high | 0.1 |
The datasets used in the FoMo-Net pre-training framework are detailed in Table 2. The proposed framework is able to process any combination of the 36 most common modalities in remote sensing, stemming from Sentinel-1, Sentinel-2, Landsat 8-9, Planet, Gaofen-2 and UAV sensors with ground sampling distance spanning from a few centimeters to per pixel. In this work, a distinct modality is defined as any band with a unique combination of wavelength and spatial resolution. Given its importance in Earth observation, digital elevation models (DEM) are also included in our definition as a specific band. The FoMo-Net pre-training pipeline is illustrated in Figure 4 and will be detailed in the following sections. Additionally, the pseudocode of the pre-training pipeline is summarized in Appendix A.
Approach to variable spectral bands: Let be a set of datasets, and be a set of spectral bands, where where and represent respectively the height and width dimensions. At each iteration, a training batch contains a variable number of spectral bands, each sampled with respective probabilities . In our experiments, we set equal to the frequency with which the -th band occurs in (see Table 2).
Each band is tokenized into patches of size according to the transformation , so that . The tokenized input is then embedded using a linear transformation learnt during the optimization process as . We considered two projection setups: The first (noted FoMo-Net1) uses a single linear projection for all spectral bands, while the second (denoted FoMo-Netm) projects each spectral band with its own linear projection. Our initial experiments suggest that FoMo-Net1 yields significantly better results.
The input to our model, then, is provided as:
where , are respectively the learnt spectral and positional embeddings, and denotes the element-wise sum.
Spectral band MAE: The next step of the procedure follows a masked autoencoding (MAE) framework [25] that is shared by various high-performing foundation models for remote sensing [60, 8, 52, 67, 16, 62, 28, 23]. However, our method differs from a typical MAE in two key aspects (see Fig. 4). First, we introduce a random spectral band selection by sampling in both and , representing the datasets and spectral bands, respectively. This approach promotes modeling interactions between a wide range of band combinations, creating a highly flexible encoder able to process any set of spectral bands. Next, we incorporate both spectral and positional embeddings, allowing the model to capture information regarding the spatial position of the patch and its spectral band, as detailed in Eq. 1. Both types of embeddings are trainable and randomly initialized. Given the randomness in the input space, the spectral embedding is of utmost importance as no other information regarding the nature of the input is provided.
Batch gradient accumulation: The combination of the proposed modality sampling and token masking is particularly computationally efficient in contrast to the exhaustive generation and processing of tokens from the entire spectrum of information. However, randomly choosing datasets and spectral bands at each training iteration results in highly heterogeneous batches, potentially destabilizing the optimization process. To address this, the FoMo-Net pre-training process exploits gradient accumulation (see Fig. 4). Letting be a neural network parameterized by , and be a loss function, the gradient is accumulated during forward passes according to , and it is this sum that is backpropagated through .
FoMo-Net architecture: In this work, our encoder is a vision transformer (ViT) [11] with 12 layers, each with 16 attention heads, and the decoder is a ViT with 8 layers, each with 16 attention heads. We use a spatial patch size of pixels. Compared to existing foundation models, FoMo-Net is quite shallow, intended as a first step towards a generic forest-monitoring foundation model, and is designed to be lightweight since most stakeholders implementing forest-monitoring algorithms in practice have very limited computational resources. In particular, FoMo-Net1 contains 101M parameters while FoMo-Netm 110M. Both FoMo-Net versions are pre-trained for 80 epochs.
Challenges: First, FoMo-Net should model interactions between a wide range of spectral bands that could potentially be utilized in a downstream task, contrary to standard foundation models that are tailored for specific sensors. Secondly, modeling data with varying GSDs require features and patterns at different scales whereas some of them are a) underrepresented and b) lack data points with wide spatial coverage. Finally, although FoMo-Net employs a data-driven strategy with limited modeling capacity since scaling up these types of architectures significantly increases the demand for computational resources.
Experimental baseline: We test the FoMo-Net pre-training framework using satellite and aerial datasets from both the FoMo-Bench and open-access satellite archives (see Table 2). We evaluate FoMo-Net on FoMo-Bench by freezing all encoder parameters and fine-tuning a task-specific head for each downstream task. We also provide an example of the potential of fully finetuning our best performing FoMo-Net on BigEarthNet-MM. Results are presented in Table 4, Table 3, and Table 5.
6 Experiments
In this section, we evaluate a range of baselines across FoMo-Bench, addressing three distinct types of computer vision tasks: classification, semantic segmentation, and object detection.
Evaluation protocols: For each task, we train a set of standard computer vision models appropriate to the task. Importantly, each of these baselines is trained in a supervised fashion on the data for that single task. This is because the heterogeneity of the tasks in both input and output data prevents the naive use of a single architecture trained on all data. (It is this challenge that we address in Section 5 with the introduction of the foundation model baseline FoMo-Net.) In the case of multi-label classification and object detection, we also evaluate a simple approach using supervised learning on multiple datasets. PolyNet denotes a ResNet-50 model pretrained on several datasets simultaneously and then finetuned and evaluated either in-domain (denoted PolyNet), i.e. on a dataset included in pretraining, or out-of-domain (denoted PolyNet), i.e. on a dataset not included in pretraining.
During the finetuning process of in-domain, out-of-domain and FoMo-Net backbones we freeze the encoder. For classiication we use a linear evaluation protocol, whereas for object detection and semantic segmentation we train a task-specific head to learn the downstream task for each dataset independently. The FoMo-Net backbone remains the same for all experiments presented in Tables 4, 3 and 5. For all experiments, we present the average performance along with the standard deviation calculated across three independent runs using different random seeds. Detailed training settings for all experiments will be provided in Appendix B.
Semantic segmentation of grid-based data: FoMo-Bench includes seven datasets suited to semantic segmentation: Waititu, ForestNet, Spekboom FLAIR #1, FLAIR #2, Woody, and FiveBillionPixels. Our baseline models here are UNet [56], DeepLabv3 [7] and UPerNet [69]. We use a Swin Transformer as a backbone for UPerNet and a ResNet50 for the rest.
| Dataset | Modalities | #C | Method | #P (Mil.) | F1 micro | mIoU |
| Waititu [33] | Aerial RGB | 3 | UNet | 32.52 | 77.96 () | 64.0 () |
| DeepLabv3 | 26.67 | 80.81 () | 67.90 () | |||
| UPerNet | 121,30 | 75.36 () | 60.56 () | |||
[…]
| FoMo-Netm | 15.70 | 71.52 () | 55.67 () | |||
| FLAIR #1 [20, 19] | Aerial multisp. | 19 | UNet | 32.52 | 72.80 () | 57.24 () |
| DeepLabv3 | 26.68 | 74.01 () | 58.75 () | |||
| UPerNet | 121.30 | 63.89 () | 46.94 () | |||
[…]
| FoMo-Netm | 25.142 | 36.6 () | 22.43 () | |||
| FLAIR #2 [18, 19] | Aerial multisp. | 19 | UNet | 32.56 | 72.99 () | 57.50 () |
| DeepLabv3 | 26.7 | 83.97 () | 72.37 () | |||
| UPerNet | 121.33 | 76.93 () | 62.52 () | |||
[…]
| FoMo-Netm | 72.32 | 25.07 () | 14.33 () | |||
| Spekboom [17] | Aerial RGB | 2 | UNet | 32.52 | 97.70 () | 95.50 () |
| DeepLabv3 | 26.67 | 97.57 () | 95.27 () | |||
| UPerNet | 121.30 | 97.81 () | 95.72 () | |||
[…]
| FoMo-Netm | 15.70 | 85.84 () | 75.19 () | |||
| FiveBillionPixels [64] | Optical | 22 | UNet | 32,52 | 74.40 () | 59.24 () |
| DeepLabv3 | 26.68 | 73.75 () | 58.42 () | |||
| UPerNet | 121.31 | 67.43 () | 51.02 () | |||
[…]
| FoMo-Netm | 20.42 | 42.45 () | 26.94 () | |||
| ForestNet [29] | Multispec. | 5 | UNet | 32.52 | 95.26 () | 90.96 () |
| DeepLabv3 | 26.68 | 95.27 () | 90.97 () | |||
| UPerNet | 121.31 | 95.23 () | 90.90 () | |||
[…]
| FoMo-Netm | 15.70 | 95.25 () | 90.93 () | |||
| Woody [35] | Aerial RGB | 4 | UNet | 32.52 | 93.37 () | 87.57 () |
| DeepLabv3 | 26.67 | 93.30 () | 87.44 () | |||
| UPerNet | 121.30 | 93.87 () | 88.46 () | |||
[…]
| FoMo-Netm | 15.70 | 87.91 () | 78.43 () |
For UNet and DeepLabv3 architectures, we use models pretrained on ImageNet. The UPerNet is based on a Swin-Base backbone from [9]. We report the mean intersection over union (mIoU) and the F1-Score at micro level for all datasets. Results are presented in Table 3.
Multi-label classification: Due to the large ground sampling distance of sensors in remote sensing, classification tasks are usually multi-label. Given a sample image, the goal is to predict all classes that are present in that image. The multi-label classification setting is thus used for BigEarthNet-MM, Sen12MS, RapidAI4EO, TreeSat AI and TalloS for either LULC or tree species recognition. Our baselines for the classification tasks include ResNet50 [26], ViT [11] and ConvNext [40] models. We initialize all models with weights pretrained on ImageNet. We pretrain PolyNet on four classification datasets within FoMo-Bench (RapidAI4EO, TalloS, Sen12MS and BigEarthNet), with their common bands (B2, B3, B4, B5, B6, B7, B8, B8a and B11 from Sentinel-2). In-domain evaluation is performed on these datasets, while TreeSatAI is excluded from pre-training to assess the encoder’s out-of-domain generalization abilities. We report micro and macro F1 scores for all datasets in Tab. 4.
Object detection The object detection task is useful in forest monitoring to count and precisely locate trees, focusing in some cases on particular species. As it requires distinguish the canopies of different trees, aerial RGB data are particularly well suited for this task. We use the NeonTree and ReforesTree datasets to test binary and multi-class object detection respectively, testing the Faster R-CNN [54], RetinaNet [39] and YOLOS [13] models (see Table 5). We measure performance by the mean intersection over union (mIoU), the mean average precision with an threshold () and the mean average precision with a variable threshold of ().
| Dataset | Modalities | #C | Method | #P (Mil.) | F1 micro | F1 macro |
| BigEarthNet [59] | Multispec. SAR | 19 | ResNet50 | 23.58 | 77.18 () | 70.32 () |
| ViT | 87.97 | 64.42 () | 47.64 () | |||
| ConvNext | 87.60 | 77.25 () | 69.66 () | |||
| PolyNet | 0.038 | 73.51 () | 66.07 () | |||
[…]
| FoMo-Netm | 0.019 | 29.15 () | 14.13 () | |||
[…]
| Sen12MS [57] | Multispec. SAR | 17 | ResNet50 | 23.57 | 51.34 () | 32.62 () |
| ViT | 87.97 | 51.49 () | 31.77 () | |||
| ConvNext | 87.60 | 50.86 () | 30.02 () | |||
| PolyNet | 0.034 | 51.70 () | 33.54 () | |||
[…]
| FoMo-Netm | 0.017 | 11.02 () | 14.29 () | |||
| TalloS (ours) | Multispec. SAR DEM | 1160 | ResNet50 | 26.34 | 18.36 () | 0.68 ( |
| ViT | 89.20 | 8.52 () | 0.24 () | |||
| ConvNext | 88.98 | 8.62 () | 0.25 () | |||
| PolyNet | 2.79 | 5.81 () | 0.19 () | |||
[…]
| FoMo-Netm | 1.39 | 7.29 () | 0.20 () | |||
| RapidAI4EO [2] | Optical Multispec. | 24 | ResNet50 | 23.60 | 62.84 () | 46.33 () |
| ViT | 88.37 | 31.61 () | 7.5 () | |||
| ConvNext | 87.61 | 64.07 ( | 47.20 () | |||
| PolyNet | 0.053 | 59.21 () | 41.11 () | |||
[…]
| FoMo-Netm | 0.026 | 12.12 () | 2.58 () | |||
| TreeSatAI [1] | Multispec. | 15 | ResNet50 | 23.56 | 53.01 () | 30.19 () |
| ViT | 87.57 | 55.48 () | 36.63 () | |||
| ConvNext | 87.60 | 64.15 () | 43.99 () | |||
| PolyNet | 0.03 | 42.41 () | 21.41 () | |||
[…]
| FoMo-Netm | 0.015 | 0.04 () | 0.019 () |
The models include ResNet-50 and DeiT-S [65] backbones pre-trained on the COCO dataset (as shown in the Table). We also include a Faster R-CNN with a DINOv2 backbone [45] fully finetuned. The FoMo-Net backbone is frozen and the evaluation is performed with a randomly initialized Faster R-CNN head.
We pretrain PolyNet simultaneously on NeonTree and ReforesTree, following which the in-domain evaluation (PolyNet) is performed on each of these datasets. For the out-of-domain evaluation (PolyNet), we pretrain the backbone on one dataset while evaluating on the other.
Semantic segmentation of point clouds: Processing point cloud representations from LiDAR sensors aims at better understanding the geometry of trees at large scale. Classifying each point, i.e. segmenting a point cloud, can be used to distinguish trees, delineate their components (e.g. trunk, branches or leaves) or evaluate their traits (e.g. height, species or biomass). FoMo-Bench includes NeonTree and FORinstance datasets for binary and multi-class point cloud segmentation respectively. For these datasets, we use the baselines PointNet [50], PointNet++[51] and Point Transformer [71]. The performances are presented in Table 6, reported with the mean intersection over union (mIoU) and the F1-score at micro level.
7 Discussion
7.1 FoMo-Bench and TalloS
We have evaluated performance on FoMo-Bench for a range of models trained on individual datasets as well as providing initial steps towards foundation models for forest-monitoring applications.
| Dataset | Modality | #C | Method | Backbone | #P (Mil.) | mIoU | ||
| NeonTree [68] | Airborne | 1 | F. R-CNN | ResNet-50 | 41.08 | 61.49 () | 38.37 () | 13.72 () |
| RetinaNet | ResNet-50 | 36.13 | 56.27 () | 6.42 () | 2.10 () | |||
| YOLOS | DeiT-S | 30.65 | 53.36 () | 28.34 () | 12.64 () | |||
| F. R-CNN | PolyNet | 14.50 | 58.42 () | 35.93 () | 11.95 () | |||
| F. R-CNN | PolyNet | 14.50 | 56.05 () | 27.77 () | 8.70() | |||
| F. R-CNN | DINOv2 | 94.36 | 44.65 () | 1.08 () | 0.22 () | |||
[…]
| F. R-CNN | FoMo-Netm | 18.59 | 43.95 () | 0.57 () | 0.10 () | |||
| ReforesTree [53] | Aerial | 6 | F. R-CNN | ResNet-50 | 41.10 | 64.67 () | 6.83 ( | 2.91 () |
| RetinaNet | ResNet-50 | 36.21 | 65.49 () | 4.69 () | 2.13 () | |||
| YOLOS | DeiT-S | 30.65 | 62.85 () | 2.30 () | 1.33 () | |||
| F. R-CNN | PolyNet | 14.52 | 62.87 () | 8.14 () | 3.30 () | |||
| F. R-CNN | PolyNet | 14.52 | 59.47 () | 7.19 () | 2.73 () | |||
| F. R-CNN | DINOv2 | 94.38 | 46.90 () | 0.08 () | 0.02 () | |||
[…]
| F. R-CNN | FoMo-Netm | 18.60 | 46.40 () | 0.12 () | 0.02 () |
Most of the presented tasks and datasets, especially for object detection and point cloud segmentation, are highly challenging and pose promising areas for future inquiry. We find that standard computer vision models such as ConvNext and Vision Transformers perform relatively poorly on several datasets, e.g. TalloS and Sen12MS. Specifically, the TalloS training set contains more than a thousand classes (see Table 1) making a single-label classification task challenging, and the multi-label classification one even more so.
In datasets such as TalloS, tree species are distributed unequally around the globe (see Tab. 2(b)), with many classes present only in specific locations. One opportunity to improve performance on such tasks could be to incorporate tree species occurrence maps as biological prior knowledge or to learn such information from the data[42].
7.2 FoMo-Net
We evaluate the performance of FoMo-Net using two pretrained backbones, FoMo-Net1 and FoMo-Netm (see Section 5) across a range of semantic segmentation, classification, and object detection tasks For each downstream task we freeze all parameters of the encoder and train a task-specific head, unless specified otherwise. For the classification tasks, we train a linear layer on top of the frozen encoder. Similarly, for the segmentation tasks we extract the encoded tokens using FoMo-Net, upsample to the input resolution and perform the pixel based classification using a multilayer convolutional segmentation head, similar to [5]. The evaluation of object detection tasks also exploits upsampled features of FoMo-Net as inputs of a randomly initialized Faster R-CNN head. In this benchmark, we report the raw performance of our models without employing task-specific strategies such as weighted loss or oversampling techniques.
Segmentation tasks: The reported performances in Table 3 highlight that FoMo-Net pre-training provides relevant semantic information. The FoMo-Net1 offers a robust foundation model baseline competing with most specialized methods for semantic segmentation, even though its finetuning process is not tailored for dense predictions. Even more, its ability to perform well on both aerial and satellite data, despite significant differences in GSD, demonstrates FoMo-Net’s versatility across multiple scales.
| Dataset | Modalities | #C | Method | #P (Mil.) | mIoU | F1 micro |
| NeonTree [68] | Airborne | 1 | PointNet | 62.12 () | 76.63 () | |
| PointNet++ | 1.45 | 55.28 () | 71.20 () | |||
| Point Transf. | 4.59 | 57.70 () | 73.15 () | |||
| FORinstance [49] | Aerial | 7 | PointNet | 35.65 () | 52.56 () | |
| PointNet++ | 1.45 | 33.00 () | 49.57 () | |||
| Point Transf. | 4.19 | 22.97 () | 37.13 () |
Classification tasks: The results presented in Table 4 indicate that the pre-training framework helped to learn significant features for LULC and tree species multi-label classification tasks, though it underperforms supervised approaches on most tasks. Remarkably, the linear evaluation setup of FoMo-Net1, with only 2.000 trainable parameters achieves SoTA performance on TalloS. Even more, when fully finetuned, FoMo-Net1 surpasses the respective ImageNet pretrained ViT architecture (denoted as FoMo-Net1-Finetuned in Table 4). It is worth noting that FoMo-Netm performs consistently worse than FoMo-Net1 in all tasks, in stark contrast to the common practice of learning a dedicated linear transformation per sensor.
Detection tasks: As shown in Table 5, both object detection tasks are challenging. Finetuning using a frozen FoMo-Net backbone performs worse than specialized models but competes with the well-established DINOv2 model with the backbone fully finetuned.
Next Steps: We have seen that our proposed FoMo-Net foundation model for forest monitoring shows promising results on some datasets but leaves considerable room for improvement. Optimizing the pre-training scheme, as well as enhancing its architecture to process non-grid like structures, e.g. LiDAR point clouds, may be especial priorities for future work. This latter direction is achievable by generating appropriate tokens from point clouds and a respective reconstruction projection to the point cloud space. Finally, it may be useful to combine the pre-training approach in FoMo-Net with supervised approaches, for a flexible combination method that can be used under variable computational constraints and with variable amounts of labeled data.
8 Conclusion
This work presents FoMo-Bench, a multi-modal, multi-scale and multi-task benchmark, geographically distributed worldwide, providing a robust and reliable evaluation framework for future forest monitoring foundation models. To further expand the task and spatial coverage of FoMo-Bench, we introduce TalloS, an extensive globally distributed dataset encompassing multi-temporal and multi-modal samples for tree species classification. Building on these contributions, we craft a foundation model, FoMo-Net, designed for forest-monitoring applications. FoMo-Net leverages a pre-training framework designed to process a flexible range of spectral bands and be applicable for a wide range of tasks in remote sensing. Our experimental evaluation of FoMo-Net on the FoMo-Bench tasks establishes the first baseline for general-purpose forest monitoring, while identifying opportunities and challenges for this promising field.
Appendix 0.A Additional details for FoMo-Net
0.A.1 FoMo-Net pretraining pseudo-code
The pseudocode of the FoMo-Net pre-training pipeline is summarized in Algorithm 1.
Appendix 0.B Evaluation settings of FoMo-Bench
In this section, we outline the training settings employed in FoMo-Bench. By default, we adhere to the training/test splits provided by the original dataset providers unless explicitly stated otherwise. Henceforth, we will denote as supervised models, all models trained end to end in a supervised fashion. All experiments in FoMo-Bench were conducted using a single Nvidia GPU, unless specified otherwise.
0.B.1 FoMo-Net settings
We pretrain both versions of FoMo-Net, namely FoMo-Net1 and FoMo-Netm, for 80 epochs using random resized crop augmentations at 128128 pixels. We set the learning rate to and the weight decay to The training utilized 4 Nvidia A100 GPUs.
0.B.2 Multi-label classification
In this section, we will provide the training settings for our multi-label classification for each dataset. All models have been trained for 20 epochs, using the AdamW optimizer. The experiments including a linear evaluation of FoMo-Net did not use weight decay, unless specified otherwise. We pretrain PolyNet for 50 epochs until the loss reaches a plateau with a learning rate of and a weight decay of . Similar to FoMo-Net we use a cosine learning rate decay schedule.
TalloS: To create the TalloS spatial splits, we divide the world map into 3-degree cells and then randomly allocate them to the train, validation, and test sets to reduce the spatial autocorrelation. Next, to ensure maximum coverage of classes within the training set, we transfer grid cells from the validation and test sets, which contain classes absent in the training set, into the training set. We train all models with a learning rate set to .We set the weight decay to with the exception of the FoMo-Net models, where we did not use weight decay.
RapidAI4EO[2]: We train on RapidAI4EO using a custom nomenclature relevant to forest monitoring. In particular, we aggregate artificial surfaces as one class. We include the following forest-related classes:
1.
Broad-leaved forest
2.
Coniferous forest
3.
Mixed forest
4.
Natural grassland
5.
Moors and heathland
6.
Sclerophyllous vegetation
7.
Transitional woodland/shrub
8.
Sparsely vegetated areas
9.
Burnt areas
10.
Glaciers and perpetual snow
Additionally, we include the following crop type classes:
1.
Non-irrigated arable land
2.
Permanently irrigated land
3.
Rice fields
4.
Vineyards
5.
Fruit trees and berry plantations
6.
Olive groves
7.
Pastures
8.
Annual crops associated with permanent crops
9.
Complex cultivation patterns
10.
Land principally occupied by agriculture, with significant areas of natural vegetation
11.
Agro-forestry areas
Finally, we include one class for water bodies and one for wetlands. RapidAI4EO is spatially organized in directories. We select two areas, i.e. 18E, 27E, as a test set. We used a learning rate of and a weight decay of for all models, with the exception of the FoMo-Net linear evaluation where we used a learning rate of without weight decay.
BigEarthNet-MM[59]: To fully finetune FoMo-Net we train for 50 epochs using a learning rate of and set the weight decay to . For the linear evaluation we set the learning rate to with no weight decay. The training of the rest of the supervised models, including PolyNet, uses a learning rate of and weight decay of .
Sen12MS[57]: For the FoMo-Net linear evaluation, we set the learning rate to . For the supervised training we set the learning rate to and the weight decay to for all models.
0.B.3 Semantic segmentation of grid-based data
All models have been trained for 20 epochs using the AdamW optimizer and a cosine learning rate decay schedule. In the following paragraphs we will describe the training settings for each segmentation dataset. FoMo-Net is adapted to a per-pixel classification scheme by building upon the embeddings of each patch of the input. The extracted features using FoMo-Net have dimensions: , where is the embedding size, is the patch size, and are respectively the height and width dimensions. Subsequently, we upsample these representations to match the input size and feed them into an autoencoder, composed of an encoder with 2 convolutional layers and 2 deconvolutional layers. The final predictions are generated using a convolution layer. No weight decay has been used for FoMo-Net finetuning unless specified otherwise.
ForestNet[29]: We upsample the input to 160160 pixels and train all models using a learning of and a weight decay of for the supervised models (no weight decay for FoMo-Net).
FiveBillionPixels[64]: We separate the scenes that will be used for train and testing and construct 120120 pixel tiles. An upsampled version, i.e. 128128 pixels, is used to train the models with a learning rate of for all models. For the supervised models we use a weight decay of .
Waititu[33]: We designate a single scene defined as the test set. From each scene in the dataset, we generate tiles of size 224224 pixels. We finetune FoMo-Net using a learning rate of with no weight decay, while the fully supervised models with a learning rate of and weight decay of .
Spekboom[17]: Similarly, we hold out 3 scenes for the test set and generate tiles for all sets. We employ a learning rate of for all models and a weight decay of for the supervised models.
Woody[35]: In a similar fashion as with the rest of the UAV datasets i.e. Waititu, and Spekboom, we hold out 3 scenes to serve as our test sets and create tiles of 224224. We train FoMo-Net using a learning rate of , while downsampling the input image to 6464. For the supervised models we use a learning rate of and a weight decay of .
FLAIR#1[20]: We use the original splits provided by the authors. The the FoMo-Net training, we downsample the input to 128128 pixels to speed up training and reduce memory requirements. We set the learning rate to . On the other hand, for the supervised models we keep the original resolution and set the learning rate to and the weight decay to .
FLAIR#2[18]: Given that the labels of the test set were not released at the time of the development, we held out 10% of the aerial scenes from the main dataset to serve as our test set. Similarly to FLAIR#1, we downsample the input size to 6464 pixels for the training of FoMo-Net while using a learning rate of . Since FoMo-Net tokenizes each channel independently, computational complexity is proportional to the number of input channels. For the supervised models, we used a learning rate of and a weight decay of , without downsampling the input image. Except for the UPerNet method where we downsample the input to 128128 pixels and set the learning rate to and the weight decay to .
0.B.4 Object detection
The training settings of all models for the object detection tasks on the NeonTree [68] and ReforesTree [53] datasets are provided in Table 7. The compared methods have either used a stochastic gradient descent (SGD) optimizer or an AdamW optimizer [41]. We also report the intersection over union (IoU) and confidence thresholds, selected within the validation splits, and applied on the predictions over the test splits. During training, we used horizontal flip, vertical flip and Gaussian noise data augmentations. We used the released splits for NeonTree, and we propose our own spatial splits for ReforesTree which are reproducible with the released code (see associated data-loader and script to create tiles from raw data). All experiments have been performed on a single NVIDIA RTX 8000 GPU.
Since DINOv2 [45] and FoMo-Net backbones are transformer-based methods, their last hidden states have been rearranged to squared dimensions and upsampled using bilinear interpolation to the input sizes. These features are processed by an autoencoder with 2 convolutional layers and followed by 2 transposed convolutional layers. A final convolutional layer generates a 32 channels tensor to fit the Faster R-CNN heads and perform the object detection tasks.
In the presented experiments, both the FoMo-Net and PolyNet backbones are frozen during the finetuning of the Faster R-CNN heads, whereas the DINOv2 backbone is fully finetuned with its corresponding heads.
| Dataset | Modality | #C | Method | Backbone | #P (Mil.) | #Epochs | Batch size | Optimizer | LR | Decay | Weight decay | IoU thresh. | Conf. thresh. |
| NeonTree [68] | Airborne | 1 | F. R-CNN | ResNet-50 | 41.08 | 200 | 64 | SGD | Exponential | 0.3 | 0.3 | ||
| RetinaNet | ResNet-50 | 36.13 | 200 | 32 | AdamW | Exponential | 0.5 | 0.3 | |||||
| YOLOS | DeiT-S | 30.65 | 300 | 10 | AdamW | Cosine | - | 0.1 | 0.99 | ||||
| F. R-CNN | PolyNet | 14.50 | 200 | 64 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | DINOv2 | 94.36 | 400 | 16 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | FoMo-Net | 18.59 | 200 | 10 | SGD | Cosine | - | 0.1 | 0.0 | ||||
| ReforesTree [53] | Aerial | 6 | F. R-CNN | ResNet-50 | 41.10 | 200 | 64 | SGD | Exponential | 0.3 | 0.3 | ||
| RetinaNet | ResNet-50 | 36.21 | 300 | 32 | AdamW | Exponential | 0.3 | 0.2 | |||||
| YOLOS | DeiT-S | 30.65 | 300 | 10 | AdamW | Exponential | 0.1 | 0.6 | |||||
| F. R-CNN | PolyNet | 14.52 | 200 | 64 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | DINOv2 | 94.38 | 500 | 16 | SGD | Exponential | 0.1 | 0.0 | |||||
| F. R-CNN | FoMo-Net | 18.60 | 400 | 2 | SGD | Cosine | - | 0.1 | 0.0 |
| Dataset | Modalities | #C | Method | #P (Mil.) | #Epochs | Batch size | LR |
| NeonTree [68] | Airborne | 1 | PointNet | 300 | 16 | ||
| PointNet++ | 1.45 | 500 | 16 | ||||
| Point Transf. | 4.59 | 500 | 16 | ||||
| FORinstance [49] | Aerial | 7 | PointNet | 300 | 16 | ||
| PointNet++ | 1.45 | 400 | 8 | ||||
| Point Transf. | 4.49 | 400 | 2 |
0.B.5 Semantic segmentation of point clouds
The training settings of all models for the semantic segmentation tasks on point clouds applied to the NeonTree [68] and FORinstance [49] datasets are provided in Table 8. The compared methods have all used the AdamW optimizer [41], with an exponential decay of and a random sampling of 1024 points per sample. During training, we used random Jitter and random rotations on the x, y, z axis as data augmentation. Note that the PointNet++ method only used random Jitter and random rotations on the z axis. We used the released splits for both NeonTree and FORinstance datasets. Both datasets provide point clouds with one million points or more per file; thus they have been spatially split in voxels containing 100,000 points at most (see associated script to create sub-point clouds from raw data).
Appendix 0.C TalloS additional information
0.C.1 TalloS - ERA5
We pair TalloS samples with the following variables from ERA5:
1.
volumetric soil water, layers 1-4
2.
skin reservoir content
3.
skin temperature
4.
soil temperature, levels 1-4
5.
surface pressure
6.
total precipitation
7.
2m temperature
8.
2m dewpoint temperature
9.
surface net solar radiation
10.
surface net thermal radiation
11.
surface solar radiation downwards
12.
surface thermal radiation downwards